According to the US Census, there are about 117,000,000 households in the US. If every household replaced 1 60W incandescent bulb that is used for 35 hours per week with a 10W LED bulb, how many kWh of electricity would this save the country per year? Based on your electric rate and your answer above, how much money would this save the country per year?
Answer #1

Answer)

Energy utilised by each household per incandescent bulb per week = 160 × 35 = 5600 Wh or 5.6 Kwh

Energy utilised per year by one household = 5.6 × 52 (as there are 52 weeks in a year) = 291.2 Kwh

Energy utilised by 117,000,000 household = 117,000,000 × 291.2 = 34,070,400,000 Kwh

Now, when 160 w bulb is replaced with 10 W led than

Energy utilised per week by one household = 35 × 10 = 350 or 0.35 kwh

Energy utilised per year per household = 0.35 × 52 = 18.2 kwh

Energy utilised by 117,000,000 household = 18.2 × 117,000,000 = 2,129,400,000 Kwh

Kwh of energy saved when shifted from 160 W bulb to 10 W bulb = 34,070,400,000 - 2,129,400,000 = 31,941,000,000 kwh

Now if we assume that each unit or each Kwh cost 8 cents than amount saves = 31,941,000,000 × 8 = 255,528,000,000 cents or 2,555,280,000 dollars.

So, it will save the country 31,941,000,000 Kwh of energy and 2,555,280,000 dollars.

Hope this helps you !!!